OpenStudy (anonymous):

Find the x-, y, and z-intercepts for the equation. Enter your answers separated by commas; example: -2, 3, 4. 2x - 6y - 3z = 12

OpenStudy (anonymous):

for x intercept, set y and z to 0, for y-intercept set x and z to 0 and so on..

OpenStudy (anonymous):

x intercept: 2x-0-0 = 12 x = 6

OpenStudy (anonymous):

you got the rest?

OpenStudy (anonymous):

I think so. Next would be 0-6*2-0=12 Making y=2?

OpenStudy (anonymous):

y intercept: -6y = 12 y = -2

OpenStudy (anonymous):

don't forget the negative sign in front of 6!

OpenStudy (anonymous):

and z is...?

OpenStudy (anonymous):

I thought it was subtraction, not a negative.

OpenStudy (anonymous):

Z-4?

OpenStudy (anonymous):

z=4 *

OpenStudy (anonymous):

2x - 6y - 3z = 12 2(0) - 6(0) - 3z = 12 -3z = 12 z = -4